OpenStudy (anonymous):

Convert 210° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

its 7pi/5 correct?

OpenStudy (anonymous):

180 degree =pi radian

OpenStudy (anonymous):

1 degree=pi/180 radian 210 degree = 210*pi/180 radian=7pi/6

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ya i was off by a lil